OpenStudy (readergirl12):

Will give free cookies..Halp :3

OpenStudy (anonymous):

post the question

OpenStudy (anonymous):

This looks like a job for @Thebadyboy

OpenStudy (thebadyboy):

i think its 34.8

OpenStudy (thebadyboy):

P=2(l+w)=2·(15.3+2.1)=34.8

OpenStudy (thebadyboy):

i just blew my own mind

OpenStudy (readergirl12):

Both of them? :o

OpenStudy (anonymous):

Mind = Blown @Thebadyboy

OpenStudy (anonymous):

I want to help you next!! These guys always beat me to it ;(

OpenStudy (readergirl12):

Wait nvm xD

OpenStudy (thebadyboy):

correct perfect equation for mind blown @Aureyliant

OpenStudy (readergirl12):

I promised more cookiez >:3

OpenStudy (anonymous):

were just cool like that @Aureyliant

OpenStudy (readergirl12):

Now eat them all till you get fatty.. xD

OpenStudy (thebadyboy):

YES!!! but i like medals BUT watever

OpenStudy (anonymous):

LOL thx @readergirl12

OpenStudy (readergirl12):

I'm about to submit it..*drum roll*

OpenStudy (readergirl12):

80% xD

OpenStudy (thebadyboy):

YEAH!!!!!!!!!!!!!!!

OpenStudy (anonymous):

I consider anything below a 96 a downright failure.

OpenStudy (readergirl12):

o-o

OpenStudy (anonymous):

Good Job @readergirl12

OpenStudy (thebadyboy):

KILLJOY @Aureyliant

OpenStudy (anonymous):

I cry I cry

OpenStudy (readergirl12):

Anyone know anything about probability? :I

OpenStudy (anonymous):

I hate probability with a passion

OpenStudy (anonymous):

But I excel in it haha

OpenStudy (readergirl12):

:o

OpenStudy (thebadyboy):

yeah me but i suck at it

OpenStudy (thebadyboy):

peace i got to fly

OpenStudy (anonymous):

Peace i gotta fly

OpenStudy (readergirl12):

Bia x3